OpenStudy (idkwut):

Let f(x) = integral e^(t^2) dt from -2 to x^2 - 3x. At what value of x is f(x) a minimum?

OpenStudy (kc_kennylau):

extremum occurs when f'(x)=0

OpenStudy (kc_kennylau):

But you should evaluate the integral first

OpenStudy (idkwut):

I got lost with all the exponents lol.

OpenStudy (kc_kennylau):

Try integral by part

OpenStudy (sidsiddhartha):

u have to use leibniz's rule

OpenStudy (kc_kennylau):

Well I can do it without the leib- whatever rule

OpenStudy (kc_kennylau):

how did the variables disappear :O

OpenStudy (sidsiddhartha):

it'll be easy with it

OpenStudy (idkwut):

1/2sqrt(pi)e(t) ? Lol, idk ,_,

OpenStudy (kc_kennylau):

From where did the pi emerge :O

OpenStudy (sidsiddhartha):

take a look http://en.wikipedia.org/wiki/Leibniz_integral_rule

ganeshie8 (ganeshie8):

\[\dfrac{d}{dx}\int \limits_{-2}^{x^2-3x }e^{(t^2)}dt = 0\]

ganeshie8 (ganeshie8):

use FTC+chain rule

ganeshie8 (ganeshie8):

\[\implies \dfrac{d}{dx}\left[f(x^2-3x) - f(-2)\right]= 0\] \[\implies f'(x^2-3x)(x^2-3x)' -0= 0\] \[\implies e^{(x^2-3x)^2}(2x-3) = 0\] \[\implies 2x-3 = 0\]

OpenStudy (idkwut):

3/2